# Emergent Hierarchy in Localized States of Organic Quantum Chains

### v1

@@ 3.2 Coupled Corrals to OQCs
Figure 6: (a) Schematics of the scattering system. The gray shaded regions correspond to left and right leads connected to the scattering region shaded in blue. (b) Conductance for the uncoupled (blue curve) and coupled (black lines) systems. Vertical dashed gray lines mark the localized states of the molecule (the ring and one precursor). Local transmission for (c) Conductive, (d) Constricted, and (e) Resonant cases, calculated at energies marked by red dots in the corresponding insets. Insets: conductance for the uncoupled and coupled systems and localized states of the molecule in dashed vertical lines.
@@
Setting the scattering problem as shown in Fig. 6(a), the electronic properties are obtained by considering the left and right leads composed of $N_{1}$-OQCs, and the center region formed by a piece of $N_{1}$-OQC attached to the molecule. The localized states of the molecule are represented by gray dashed vertical lines exhibited together with the conductance results for the uncoupled N1-OQC and the coupled system composed of chain+precursor+corral, represented by blue and black curves, respectively, in Fig. 6(b). The attachment of the molecule to the OQC affects the ballistic transport of the original $N_{1}$-OQC (blue curves) usually reducing the typical 1D channel of the $2e^{2}/h$ quantum flux. However, at particular energy values, more dramatic changes are achieved, such as full conduction suppression or resonant state.
@@
When spatially coupled, molecular states and chain continuum states may coexist, giving rise to different responses, here distinguished by conductive, constricted and resonant cases, as shown in the normalized local transmission results in Fig. 6(c-e) with insets highlighting each particular physical phenomena. The local transmission results of panels (c) and (e) are examples in which the $G(E)$ values match perfectly with energy states of the molecule (red dots). However, in the resonant case, the electrons are trapped at very narrowed energy ranges, exhibiting almost ballistic transport. For the conductive case, the molecular states seem to be absorbed by the continuum of the OQC. Moreover, such local transmission exhibits a preferential central route that crosses the middle of the chain, although all bonds of the chain are effectively visited, differently from those of the corral sites that do not contribute to transmission. Interestingly, the resonant case is characterized by a local transmission that occurs in both chain and molecule, due to a small but sufficient percolation wave probability inside the precursor connecting chain and corral (gray shades). In fact, the sign is amplified at the internal part of the coupled corral indicated by purple in the color bar, as shown in Fig. 6(e).
@@
Generally speaking, several localized states of the molecule have trapped waves into the continuous band of the coupled system (molecule+chain). In the coupled regime, emergent resonant states of the finite system ($E_{res}$) may lead to highly localized coupled modes within the continuum states, resulting in bound states in the continuum (BIC). To measure the intensity of this sign in electronic devices, we define a dimensionless quantity, $\mathcal{Q}=\frac{|E_{res}|}{\Delta E_{res}}$, where the strength of this interaction is related to the Fano antiresonance width $\Delta E_{res}$. The $\mathcal{Q}$ factor provides a measure of the quality of the resonant state. It directly quantifies spectral broadening and hybridization, indicating the extent to which the sharp, discrete energy levels of the isolated molecule are absorbed into the continuous spectrum of the chain upon coupling. By definition, true BICs are related to $\Delta E_{res}\rightarrow 0$, and consequently $\mathcal{Q}\rightarrow\infty$ [47]. Considering the quality factor for the present resonant example [Fig. 6(e)], we notice that the coupled state, in the black curve, almost fulfills the condition $\Delta E_{res}\rightarrow 0$, being defined as a quasi-BIC state. The molecular state, in gray, is responsible for the resonance peak (black curve) corresponding to $G(E)$ close to one quantum conductance flux, similarly to the continuum states of the chain (blue lines).
@@
For some energies of the injected current, zero conductive is provided by electronic destructive interference. As shown in the constricted case in Fig. 6(d), the state does not fully percolate through the central part, being spatially trapped between the left lead and the molecule.
4 Conclusions
@@
In this work, we have presented a theoretical analysis of experimentally synthesized organic quantum chains (OQCs), combining density functional theory (DFT) and molecular dynamics (MD) to obtain relaxed structures. The electronic properties were accurately captured by a tight-binding (TB) model with an exponential hopping function within the ab initio relaxed structures, being in good agreement with DFT calculations.
@@
A key consequence of the quasi-one-dimensional character of these systems is the formation of singular hierarchical conductance channels within the electronic spectra. The confinement is enhanced by inserting multiple precursors within the unit cell, leading to a partitioning of continuous states, creating a hierarchy of pseudo-gaps. This electronic structure makes these materials highly promising for multi-band light-emitting devices, as it promotes controlled electronic transitions between the emergent sub-bands. Furthermore, the molecular chains, are characterized by an scale invariant large energy gap of $\Delta E=2.0$ eV for all proposed $N_{i}$ OQCs. This prediction matches very well with the measured experimental value [1].
@@
Our analysis also revealed distinct localization phenomena. Compact localized states (CLS) were identified at fundamental energy levels, where the probability density is confined to a small number of sites within the unit cell. Additionally, other robust localized states, manifested as very narrow conductance channels, were observed to persist regardless the increasing number of precursors. Finally, we propose a quantum device based on the experimental formations of the carbon structures. The system composed of a OQC attached to a quantum corral shows prominent interferences and bound states in the continuum in the conductance results. For conductive situations, the transmission probability occurs in the very center of the OQC, being completely suppressed along the coupled corral. Resonant cases appear as quasi-BICs with transmission occurring along both chain and molecule, being amplified inside the corral.
@@ Acknowledgements
Appendix A Structure and Energy of OQCs
With respect to bond lengths obtained by the methods described in Section II, all $\text{N}_{i}$-OQC systems display similar structural features regardless of the chain length. The systems consist exclusively of carbon and hydrogen atoms forming sp2 hybridizations with five- and six-fold carbon rings, the edges of which are passivated by hydrogen atoms. In pentagonal rings, the bond lengths C – C are approximately 1.47 Å. In hexagons, these lengths range from 1.39 Å to 1.43 Å. The inter-unit C – C bonds connecting the molecular precursors measure approximately 1.49 Å. The C – H bonds exhibit the typical value of 1.10 Å, and are not included in TB computations.
Table A1: Unit cell length $l_{z}$ for Ni-OQC systems.
[…]
N1 $19.4$
N2 $37.4$
N3 $54.8$
N4 $65.3$
N5 $72.2$
N6 $73.0$
N7 $68.8$
N8 $61.3$
N9 $61.8$
N10 $50.4$
More details on the length of the unit cell are shown in Table A1, where an increase in size is observed for i=16i=1-6, and a decrease for $i>6$ chains. This result underscores the relationship between chain-type and corral-type topologies. The argument of enhanced stability for the $N_{12}$ organic quantum corral proposed by Peng et al. [1] is fundamentally linked to its geometry. In the case of the open OQC structures considered here, the threshold at $i=6$ aligns with the corral configuration, as the systems begin to expand laterally and display an increased curvature beyond this point. Nearly identical $l_{z}$ values are observed for $i=8,9$, followed by a significant reduction in $l_{z}$ for $i=10$, when inter-chain interactions become significant.
From an energetic perspective, the cohesive energy of the $\text{N}_{i}$-OQC systems was calculated according to the expression,
$E_{\text{coh}}(\text{N}_{i}\text{-OQC})=\frac{E_{\text{T}}(\text{N}_{i}\text{-OQC})-40iE_{\text{C}}-20iE_{\text{H}}}{60i}$ (8)
where $E_{\text{T}}(\text{N}_{i}\text{-OQC})$ is the total energy of the $\text{N}_{i}$-OQC configuration, and $E_{\text{C}}$ and $E_{\text{H}}$ are the total energies of the isolated carbon and hydrogen atoms, respectively. This formulation normalizes the total energy per atom, enabling a direct comparison across different system sizes and topologies.
For all configurations investigated in this work, in contrast to the circular topologies reported in Ref. [1], the cohesive energy remains constant, with $E_{\text{coh}}(\text{N}_{i}\text{-OQC})=\text{-7.99}$ eV/atom, regardless of the value of $i$. In the case of quantum corrals, the minimal cohesive energy was reported for $N_{12}$, which corresponds to $N_{6}$-OQC in the present notation. This difference arises from the topological distinctions between open-chain and closed-loop geometries. The open-chain structures studied here are not subjected to the same geometric constraints that force the molecules into less favorable conformations because the chains are free to adapt their curvature while maintaining energetically equivalent configurations. Furthermore, our investigation extends up to $i=10$, which corresponds to the limit where chain-chain begin to occur, via H-H interactions, within the OQC structure, as illustrated in Fig. 1(m). For values of $i$ greater than ten, additional topological constraints are expected to impact the stability and morphology of the system.

### v2

@@ III.2 Coupled Corrals to OQCs
Figure 5: (a) Schematics of the scattering system. The gray shaded regions correspond to left and right leads connected to the scattering region shaded in blue. (b) Conductance for the uncoupled (blue curve) and coupled (black lines) systems. Vertical dashed gray lines mark the localized states of the molecule (the ring and one precursor). Local transmission for (c) Conductive, (d) Suppressed, and (e) Resonant cases, calculated at energies marked by red dots in the corresponding insets. Insets: conductance for the uncoupled and coupled systems and localized states of the molecule in dashed vertical lines.
@@
The transport properties of the coupled system are explored by setting the scattering problem as shown in Fig. 5(a), i.e. considering the left and right leads composed of $N_{1}$-OQC, and the center region formed by a piece of $N_{1}$-OQC attached to the molecule. Localized states of the molecule are represented by gray dashed vertical lines exhibited together with the conductance results for the uncoupled N1-OQC and the coupled system composed of chain+precursor+corral, represented by blue and black curves, respectively, in Fig. 5(b). The molecule coupling to the OQC affects the ballistic transport of the original $N_{1}$-OQC (blue curves) usually reducing the typical 1D channel of the $2e^{2}/h$ quantum conductance. However, at particular energy values, more dramatic changes are achieved, such as full conduction suppression or resonant state.
@@
When spatially coupled, molecular states and the continuum states of the chain can coexist, giving rise to distinct transport responses, here classified as conductive, suppressed, and resonant, as illustrated by the normalized local transmission maps in Fig. 5(c-e) with the insets highlighting the corresponding physical regimes. The local transmission patterns in panels (c) and (e) correspond to cases in which the local transmission is calculated exactly at a discrete molecular energy level (red dots in the conductance plot). In the resonant regime, electrons are confined to very narrow energy windows, exhibiting almost ballistic transport. In contrast, in the conductive regime, the molecular states are effectively absorbed into the continuum of the OQC. In the latter case, the local transmission displays a preferential central route along the chain, although all chain bonds are effectively visited, in contrast to the corral sites, which do not contribute to transmission. Interestingly, in the resonant case the local transmission extends over both the chain and the molecule, owing to a small but finite percolation of the electronic states through the precursor connecting the chain and the corral (gray shades). In fact, this contribution is further enhanced inside the coupled corral region, as indicated by the purple color scale in Fig. 5(e).
@@
In general, several molecular localized states become embedded in the continuous of the coupled system (molecule+chain). In this coupled regime, emergent resonant states of the finite system ($E_{res}$) can give rise to strongly localized hybrid modes within the continuum, resulting in the formation of bound states in the continuum (BIC). The intensity of this effect in electronic devices, may be quantified by a dimensionless quality factor, $\mathcal{Q}=\frac{|E_{res}|}{\Delta E_{res}}$, where $\Delta E_{\mathrm{res}}$ denotes the width of the Fano antiresonance and characterizes the coupling-induced spectral broadening. The factor $\mathcal{Q}$ provides a measure of the quality of the resonant state, quantifying the degree of hybridization and the extent to which the sharp, discrete energy levels of the isolated molecule are absorbed into the continuum spectrum of the chain upon coupling. By definition, true BICs [53] correspond to the limit $\Delta E_{res}\rightarrow 0$, and consequently $\mathcal{Q}\rightarrow\infty$. For the resonant example shown in [Fig. 5(e)] the calculated quality factor indicates that the coupled state (inset black curve) nearly satisfies the condition $\Delta E_{res}\rightarrow 0$, being defined as a quasi-BIC state. The molecular state (gray dashed line) gives rise to the resonance peak in the conductance, with $G(E)$ approaching to one quantum conductance flux, in close analogy with the continuum states of the chain (blue lines).
@@
For some energies of the injected current, zero conductance is achieved by the electronic states, due to destructive interferences that may occur inside the scattering region. As shown in the suppressed case in Fig. 5(d), the state does not fully percolate through the central part, being spatially trapped between the left lead and the molecule.
Conclusions
@@
In this work, we have presented a theoretical analysis of experimentally synthesized organic quantum chains, combining DFT and CMD to obtain relaxed structures. The electronic properties were accurately captured by a TB model with an exponential hopping function within the ab initio relaxed structures, being in good agreement with DFT calculations.
@@
A key consequence of the quasi-one-dimensional character of these systems is the formation of singular hierarchical conductance channels within the electronic spectra. The confinement is enhanced by inserting multiple precursors within the unit cell, leading to a partitioning of continuous states and a densification of narrow conductance plateaus, leading to a pseudo-band hierarchy. Our analysis revealed an interesting localization phenomenon governed by the molecular nature of the OQC building blocks. We identified confined molecular states near the Fermi level associated with electronic carriers trapped within specific precursor sub-units due to weak inter-unit coupling. These states appear as flat bands in the electronic spectrum and do not contribute to ballistic transport. This electronic structure makes these materials highly promising for multi-band light-emitting devices, as it promotes controlled electronic transitions between the emergent sub-bands. Furthermore, the molecular chains, are characterized by an scale invariant large energy gap of $\Delta E\approx 2.0$ eV for all proposed $N_{i}$ OQCs. This prediction matches very well with the measured experimental value [1].
@@
Finally, we propose a quantum device based on the experimental formations of the carbon structures. The system composed of a OQC attached to a quantum corral shows prominent interferences and bound states in the continuum in the conductance results. For conductive situations, the maximum transmission probability occurs in the very center of the OQC, being completely suppressed along the coupled corral. Resonant cases appear as quasi-BICs with transmission occurring along both chain and molecule, being amplified inside the corral.
@@
Author contributions
All authors contributed to the design of the research, the analysis of the results, and the writing of the manuscript. L.L.L. and A.B.F. performed the numerical calculations.
Conflicts of interest
There are no conflicts to declare.
Data availability
The code and files used within Tight-binding calculations can be found at https://github.com/lucaslopeslage/OQC_code_and_files.git.